OpenStudy (anonymous):

Find an equation of the parabola of the form y=ax^2+bx+c (1,6),(-1,8),(3,36) ???

OpenStudy (queelius):

System of equations. Three unknowns, three equations. We are given that the given three points all fall on a parabola of the form y = ax^2 + bx + c. So, we have to figure out which values for a, b, and c will satisfy those given points. Using the points given, we know that: 6 = a + b + c 8 = a - b + c 36 = 9a - 3b + c So, solve the system of equations and you have your answer.

OpenStudy (anonymous):

a=13/4, b=-1, c=15/4 this is what i got but it was not correct after in entered it in the equation

OpenStudy (queelius):

I made an error. The last equation should be 36 = 9a + 3b + c. Try it again.

OpenStudy (anonymous):

Thank you